amy travels 450 miles at a certain speed. If the car had gone 15 mph faster, the trip would have take 1 hour less. Find amys speed. Show your work.
I know that D=RT So 450=RT I am not sure where to go from here because this does not work
450=(15+r)(t-1) ---- any suggestions? I appreciate the help, thank you

Answer by richwmiller(17219)

There is nothing wrong with your equation
450=(r+15)*(t-1)
You forgot the original info.
r*t=450
r=75 and t=6
check
r*t=450
75*6=450
450=450
now check the second equation
90*5=450
450=450
ok
The trip would not have been one hour less because Amy would have gotten stopped and ticketed or most likely ended up as a traffic fatality (dead).
What a terrible problem for math!